3. The demand and supply for hospital care are given by the equations

Q = 400 – P and Q = 80 + 3P, respectively.

A. Solve for the equilibrium quantity and price. (5 pts).

B. Now suppose that patients have insurance so that they only pay 20% of the price for hospital care. Solve for the new equilibrium quantity and price. (5 pts)

4. Consider the following demand and supply data

Price   Demand            Supply

0         200                  0

1         90                    50

2         80                    70

3         60                    60

4         40                    100

5         0                      150

  1. Draw the supply and demand curves using a graph (5 pts)

B. Describe the situation if Price = 4 (5 pts)
Answer #1

Q3) at eqm

Qd = Qs

400-P = 80+3P

320 = 4P

P* = 80, Q* = 80+3*80 = 320

(P,Q) = (80,320)

B) new demand curve

Q = 400- .2P

At new eqm, 400-.2P = 80+3P

320 = 3.2P

P' = 100, Q' = 80+3*100 = 380

(P,Q)' = (100,380)

Q4)

Graph

B) p = 4

Qd = 40, Qs = 100

As Qd < Qs, so excess supply, surplus creates in Economy